# The Physics and Metaphysics of Social Powers: Bridging Cognitive Processing and Social Dynamics, a New Perspective on Power through Active Inference

### v1

@@ 0.2 Attention as power: shared narratives and awareness or coercion
Power as potential, as force: as the ability to do work in physics and the ability to do work in the socioreproductive realm, represent two sides of the same coin: the mechanics of matter and energy sustain the dynamics of complex systems; which, in turn, modulate matter and energy. Under AIF, this ‘common currency’ can be studied as the process by which state spaces are explored by collectives (dis)attending to projective possibilities. In order to understand this better, formalizations are necessary, which is our interest in this paper: awareness about the modulation of attention begins at the level of understanding the constraints framing attention. As Foucault notes in “The Subject and Power” (1982), formalizations are not without their reductive, objectifying problems, but are necessary in order to drive critical analysis forward. Further, he also notes: “What we need is a new economy of power relations—the word “economy” being used in its theoretical and practical sense” (p. 779). This is, in great part, what we aim to explore here, not an established structure but “a technique, a form of power” (p. 781): attention itself, analyzed under AIF.
@@
In the socioconstructive realm we currently observe: those under duress resort to inevitably short-term policies, as well as pool their attentional trust in those who seem to promise them less uncertainty131313“…in trials in which the agents’ prior beliefs are strong, such as after a change of mind, they do not simply follow one another, but try to fulfill their prior belief—and this explains why we observe several errors after only one of the agents changes mind. These examples illustrate that it is the strength (or the precision) of the beliefs about the joint goal context that determines whether or not an imitative response takes place.” Maisto et al, 2024.. Those who can rely on the minimization of uncertainty because of historically-given, infrastructural factors have access to more and therefore longer-term leverage on the unfolding future. Power as the exertion of longer-term agency implies that, to begin with, an agent relies on factors that make more spacetime available for the computation (as projection: planning and effectuation) of possible futures.
@@ 1 Theoretical Framework: Active Inference
Interestingly, we can also link it to an enactive inference perspective, which focuses more strongly on the role of action in shaping perception. Under this lens, beliefs are formed actively. For readers unfamiliar with the intricacies of Active Inference Framework (AIF), this concept can be seen as a dynamic process where agents optimize their interactions with and models of the world to reduce uncertainty and maintain homeostasis, allostasis, or homeorhesis (Ororbia and Friston, 2023). AIF highlights the importance of generative models that represent the structure of the environment and guide behavior. It directly integrates perceptual, cognitive, and motor processes.
Generative models are thus the heart of any AIF agent. They represent a structured way for agents to analyze sensory inputs and infer their possible causes. These models consist of hierarchical representations that generate predictions about sensory states based on beliefs (i.e., inductive habits) about hidden causes. Generative processes represent the true states of the world that produce sensory observations, while generative models represent the agent’s internal hypotheses of these processes. AIF hinges on the alignment of models with processes, with actions taken to correct any mismatch. This framework allows agents to learn both the structure of the environment and effective policies through message-passing algorithms, which mimic neuronal processes (Parr et al., 2019).
@@
In generative models, the primary components involved in sensemaking include observable sensory states, hidden states, and parameters that encode the probabilistic dependencies between them: Sensory states are the inputs that the agent receives from the environment. They are the observations the model is trying to predict or explain based on hidden states and other parameters. Sensory states provide feedback to the generative model, helping the agent assess how well its predictions align with real-world data. This can be represented as $P(o\mid s,\theta)$, where $s$ is a hidden state, and $\theta$ represents the parameters of the generative model. Hidden states capture the unobserved causes or latent variables that the agent infers to explain sensory states. Hidden states can represent underlying factors or situations that produce the sensory input, which aren’t directly measurable but can be inferred from patterns in sensory states. In Bayesian terms, these are inferred by the model to reduce uncertainty (or surprise) about observable data. The distribution over hidden states can be represented as $P(s\mid\theta)$, where $\theta$ (i.e., the generative model parametrization) influences the hidden state predictions based on prior beliefs or learned parameters. Parameters define the relationships between hidden and sensory states. Parameters capture the structural and statistical dependencies within the model, and they are updated over time to maximize model fitness to the data. Parameters specify the exact form of dependencies, such as the strength or variance of the connection between hidden states and observable outcomes.
@@ 1.1 Free energy minimization
Free energy minimization is the unifying principle that motivates perception, cognition, and action in Active Inference Framework (AIF). As described in Parr and Friston’s 2019 paper, this minimization occurs at both local and larger scales. Free energy minimization at cellular and cognitive levels could reflect larger systemic constraints and efficiencies. This process involves reducing the divergence between an agent’s predictions and its sensory inputs, maintaining a state of low surprise or entropy.
Free energy is an upper bound on the “surprise” (or negative log-probability) an agent experiences given its observations ($o$). Surprise reflects the degree to which sensory data diverges from an agent’s predictions, but it is generally intractable to compute directly. Variational free energy (VFE), an approximation applied by Friston and others to the study of the brain (Friston et al., 2006, Friston et al., 2016, Friston et al., 2015), allows the agent to sidestep this intractability by using an internal generative model, represented as $q$-distributions over hidden (latent) states ($s$), to approximate the posterior distribution. The variational free energy $F$ can be expressed as:
@@
$\mathbb{E}_{q(s)}[\ln p(o\mid s)]$ represents the expected log-likelihood of observations given the hidden states. Maximizing this term makes the agent’s model predictive of actual sensory inputs. To minimize free energy, and align their sensory inputs with their predictions, agents have to update their beliefs about the world. Belief updating under AIF is achieved through variational Bayesian inference, whereby agents continuously update their beliefs about the state of the world based on new sensory information. This allows agents to make predictions and adjust their actions to confirm these predictions, thereby preserving a coherent sense of self and environment.
@@
Minimization can be broken down into two complementary strategies. Agents can perform perceptual or active inference, which are inextricably related. Perceptual inference entails adjusting beliefs about hidden states to make them more consistent with current sensory data, which is akin to learning or updating internal representations to match the environment. AIF, on the other hand, entails modifying the environment or seeking new sensory data to make incoming data consistent with beliefs, guiding action selection and resulting in a feedback loop where actions bring sensory data closer to expectations. It is important to note difficulty in differentiation between the active and the perceptive, which brings us to the question of modeling collective desire: is there a way to tell the difference between when we act according to an “inner” logic we have conscious cognitive access to, and when are we moved by a logic we simply ‘feel’ driven by? Parr and Friston’s work (2019) suggests that free energy minimization does not operate only at one scale (e.g., neuronal or cognitive levels) but may extend across multiple biological scales, from cellular to social systems. Friston and others (Vasil et al, 2020) postulate that free energy minimization, as a principle, might reflect broader systemic efficiencies in biological scaling laws, where living systems seek to maintain optimal internal states (allostatic setpoints) with minimal energetic cost. The “average” aspect is important to emphasize here: for complex systems long-term averages may signify significant shorter-term high-risk, high-energy investments.
@@
where $\ln\tilde{p}(o)$ represents the agent’s preferences over observations or extrinsic value, expressed as a (biased) prior probability distribution over observations (Parr et al., 2022). The first term represents the (negative) expected information gain or epistemic value, which scores the degree to which an agent expects to update their beliefs in light of new sensory data. By pursuing outcomes that maximize a balance between epistemic and extrinsic values, agents simultaneously satisfy immediate goals and resolve uncertainty, driving adaptive behavior in complex environments.
@@ 1.2 Why active inference is suitable for modeling power dynamics
Active Inference Framework (AIF) is well-suited to understand power dynamics, beginning with individual agents as predictive systems that minimize uncertainty through perception and action (Bruineberg, 2017). These agents develop control through embodied interaction with their environment, establishing what Bruineberg calls the “primacy of the ’I Can”’—the fundamental ability, or power, to influence one’s surroundings. This individual control scales up through generative models that enable agents to predict and understand others’ intentions (Friston et al., 2020; Vasil et al., 2020), while attention is fundamentally shaped by deep preferences that determine what becomes salient in the environment (Parvizi-Wayne, 2024).
@@
At the social level, power dynamics emerge through shared predictive models and environmental mediation. Groups develop shared narratives that coordinate actions and expectations (Friston et al., 2020), while stigmergic processes allow for distributed control through environmental modifications that influence collective behavior (Friedman et al., 2021). This social coordination is reinforced through deontic cues that signal social norms and expectations (Constant et al., 2019), while cultural practices shape attention patterns to create shared perceptual landscapes within the environment. groups (Wen, 2024). Joint attention and shared fields of affordances enable efficient coordination through mutual understanding of action possibilities (Tison and Poirier, 2021; Vasil et al., 2020).
@@
These mechanisms ultimately can give rise to complex power structures through decentralized interactions rather than top-down control (Friedman et al., 2021). Groups self-organize into subgroups based on belief similarity (Kastel et al., 2023), while influence spreads through network structures that amplify certain messages while marginalizing others (Albarracin et al., 2022). This process creates systems of ‘cultural capital’ (Bourdieu) where shared knowledge and norms enable efficient navigation of social contexts (Constant et al., 2019; Hipólito and van Es, 2022). The result is a dynamic system where power emerges from the interplay of individual prediction, social influence, and cultural learning, all of which can be unified under the principles of AIF.
@@
The question of social sameness (as in “same” narratives, belief similarity, etc.) and (dis)attention to it, is therefore a political question. If attention is fundamentally shaped by deep preferences that determine what becomes salient in the environment for an agent or group, then we ought to ask how the production and performance of scripts (habitus, narratives: histories, identities, etc.) can be newly understood so that we may critically construct these collectively, without risking that this collectivity is induced by strong homogenizing scripts which occlude or incapacitate ever-renewing possibilities (as warned by Arendt). Incentives to promote legibility in a population can also manifest in institutions shaping and direction of attention toward specific narratives (Scott, 2020), and may similarly arise in collectively determined scripts. Thus, a balance needs to be maintained between benefits such as improved coordination and an oversuppression of diversity among constituent members of collectives. As mentioned earlier, “There is no alternative” can be understood as the poster-child of ‘one-dimensional’ [^14] For this Marcusian notion at work in similar ways in the domain of NLP, see de Jager 2023a. visions effectuated by reducing complex systems into simplified structures which pool attention towards a specific goal. Through our treatment, a new definition of social power can thus be understood as possibilistic. Essentially: if we ought to adopt a politics based on the knowledge that scripts work in the ways discussed above and in what follows below, and if the organization of life is hierarchical by definition (both physically and metaphysically: we are given to a world of already-existing constraints which we organize around by means of concepts aimed at future possibilities), taking account of this by understanding the nature of distributed generative models—over time: as generations, as the representation of historical narratives, etc.—offers a vision of possibilistic landscapes based on collective dynamics (and their analyzable mechanics) which can be framed beyond our current visions based on simplistic ideas of agency as individualized freedom and through atomized actors.
@@ 2.1 Scripts as frameworks for social behavior
Social scripts are blueprints that enable agents to navigate social situations and select situationally appropriate behaviors (Albarracin et al., 2021). They represent both internalized cognitive schemas and externalized social orders that structure behavioral possibilities within cultural contexts. Scripts harness socially shared knowledge that prescribes normative standards for behavior, similar to how actors share a dramaturgical script. Strong scripts specify precise sequences of structured behavioral events that agents are expected to perform in certain social situations (Albarracin et al., 2021). These scripts maintain causal and temporal relationships between events, enabling agents to make reliable inferences about others’ likely actions. For example, the North American restaurant script prescribes specific ordered sequences: being seated by the host, reviewing menus, ordering drinks, ordering food, eating, and paying the bill including tip (Albarracin et al., 2021).
@@
Weak scripts specify the typical features and semantic associations that agents encounter in particular event types, without strictly prescribing event order (Albarracin et al., 2021). These scripts function as clusters of conceptually related possibilities that help agents interpret situations and identify appropriate actions. The semantic connections in weak scripts adapt agents’ perceptual fields by making certain action possibilities more salient based on contextual goals.
@@
Scripts help maintain power structures. Scripts normalize certain behavioral patterns by providing templates for ”appropriate” action, making deviations appear abnormal or inappropriate (Constant et al., 2018). This normalization reinforces existing power relations by making them appear natural and inevitable. They establish reliable sequences of interaction in favor of a current norm: whether agents following this are (critical) aware of the norm, or simply perform it, is often the question. They thus reduce uncertainty and create predictable social patterns that benefit those already in power (Albarracin et al., 2021). These patterns become self-reinforcing as agents learn to expect and reproduce them. Given this predictability, scripts assign and maintain social roles by specifying appropriate behaviors for different agents based on their position within power structures (Constant et al., 2018). This role differentiation helps reproduce hierarchical relationships.
@@
Looking at scripts under AIF, they can be understood as shared generative models that enable coordinated social behavior. Scripts represent shared knowledge about the causal structure of social situations, allowing agents to make predictions about others’ behavior and select appropriate actions. These shared models facilitate social coordination by reducing uncertainty about interaction patterns. Scripts incorporate deontic cues that trigger specific behavioral policies, helping agents select appropriate actions in response to environmental signals (Constant et al., 2018). These cues guide action selection by making certain policies more likely in specific contexts. Scripts are transmitted across generations through cultural learning processes, enabling the reproduction of social structures. This transmission occurs through both explicit instruction and implicit learning from environmental regularities.

### v2

@@ 0.2 Attention as power: shared narratives and awareness or coercion
Power as potential, as force: as the ability to do work in physics and the ability to do work in the socioreproductive realm, represent two sides of the same coin: the mechanics of matter and energy sustain the dynamics of complex systems; which, in turn, modulate matter and energy. Under AIF, this ‘common currency’ can be studied as the process by which state spaces are explored by collectives (dis)attending to projective possibilities. In order to understand this better, formalizations are necessary, which is our interest in this paper: awareness about the modulation of attention begins at the level of understanding the constraints framing attention. As Foucault notes in “The Subject and Power” [Foucault, 1982], formalizations are not without their reductive, objectifying problems, but are necessary in order to drive critical analysis forward. Further, he also notes: “What we need is a new economy of power relations—the word “economy” being used in its theoretical and practical sense” (p. 779). This is, in great part, what we aim to explore here, not an established structure but “a technique, a form of power” (p. 781): attention itself, analyzed under AIF.
@@
In the socioconstructive realm we currently observe: those under duress resort to inevitably short-term policies, as well as pool their attentional trust in those who seem to promise them less uncertainty131313“…in trials in which the agents’ prior beliefs are strong, such as after a change of mind, they do not simply follow one another, but try to fulfill their prior belief—and this explains why we observe several errors after only one of the agents changes mind. These examples illustrate that it is the strength (or the precision) of the beliefs about the joint goal context that determines whether or not an imitative response takes place.” [Maisto et al., 2024].. Those who can rely on the minimization of uncertainty because of historically-given, infrastructural factors have access to more and therefore longer-term leverage on the unfolding future. Power as the exertion of longer-term agency implies that, to begin with, an agent relies on factors that make more spacetime available for the computation (as projection: planning and effectuation) of possible futures.
@@ 1 Theoretical Framework: Active Inference
Interestingly, we can also link it to an enactive inference perspective, which focuses more strongly on the role of action in shaping perception. Under this lens, beliefs are formed actively. For readers unfamiliar with the intricacies of Active Inference Framework (AIF), this concept can be seen as a dynamic process where agents optimize their interactions with and models of the world to reduce uncertainty and maintain homeostasis, allostasis, or homeorhesis [Ororbia and Friston, 2023]. AIF highlights the importance of generative models that represent the structure of the environment and guide behavior. It directly integrates perceptual, cognitive, and motor processes.
Generative models are thus the heart of any AIF agent. They represent a structured way for agents to analyze sensory inputs and infer their possible causes. These models consist of hierarchical representations that generate predictions about sensory states based on beliefs (i.e., inductive habits) about hidden causes. Generative processes represent the true states of the world that produce sensory observations, while generative models represent the agent’s internal hypotheses of these processes. AIF hinges on the alignment of models with processes, with actions taken to correct any mismatch. This framework allows agents to learn both the structure of the environment and effective policies through message-passing algorithms, which mimic neuronal processes [Parr et al., 2019].
@@
In generative models, the primary components involved in sense-making include observable sensory states, hidden states, and parameters that encode the probabilistic dependencies between them: Sensory states are the inputs that the agent receives from the environment. They are the observations the model is trying to predict or explain based on hidden states and other parameters. Sensory states provide feedback to the generative model, helping the agent assess how well its predictions align with real-world data. This can be represented as $P(o\mid s,\theta)$, where $s$ is a hidden state, and $\theta$ represents the parameters of the generative model. Hidden states capture the unobserved causes or latent variables that the agent infers to explain sensory states. Hidden states can represent underlying factors or situations that produce the sensory input, which aren’t directly measurable but can be inferred from patterns in sensory states. In Bayesian terms, these are inferred by the model to reduce uncertainty (or surprise) about observable data. The distribution over hidden states can be represented as $P(s\mid\theta)$, where $\theta$ (i.e., the generative model parametrization) influences the hidden state predictions based on prior beliefs or learned parameters. Parameters define the relationships between hidden and sensory states. Parameters capture the structural and statistical dependencies within the model and are updated over time to maximize model fitness to the data. Parameters specify the exact form of dependencies, such as the strength or variance of the connection between hidden states and observable outcomes.
@@ 1.1 Free energy minimization
Free energy minimization is the unifying principle that motivates perception, cognition, and action in Active Inference Framework (AIF). As described in Parr and Friston’s 2019 paper, this minimization occurs at both local and larger scales [Parr et al., 2019]. Free energy minimization at cellular and cognitive levels could reflect larger systemic constraints and efficiencies. This process involves reducing the divergence between an agent’s predictions and its sensory inputs, maintaining a state of low surprise or entropy.
Free energy is an upper bound on the “surprise” (or negative log-probability) an agent experiences given its observations ($o$). Surprise reflects the degree to which sensory data diverges from an agent’s predictions, but it is generally intractable to compute directly. Variational free energy (VFE), an approximation applied by Friston and others to the study of the brain [Friston et al., 2006, 2015, 2016], allows the agent to sidestep this intractability by using an internal generative model, represented as $q$-distributions over hidden (latent) states ($s$), to approximate the posterior distribution. The variational free energy $F$ can be expressed as:
@@
$\mathbb{E}_{q(s)}[\ln p(o\mid s)]$ represents the expected log-likelihood of observations given the hidden states. Maximizing this term makes the agent’s model predictive of actual sensory inputs. To minimize free energy and align their sensory inputs with their predictions, agents have to update their beliefs about the world. Belief updating under AIF is achieved through variational Bayesian inference, whereby agents continuously update their beliefs about the state of the world based on new sensory information. This allows agents to make predictions and adjust their actions to confirm these predictions, thereby preserving a coherent sense of self and environment.
@@
Minimization can be broken down into two complementary strategies. Agents can perform perceptual or active inference, which are inextricably related. Perceptual inference entails adjusting beliefs about hidden states to make them more consistent with current sensory data, which is akin to learning or updating internal representations to match the environment. AIF, on the other hand, entails modifying the environment or seeking new sensory data to make incoming data consistent with beliefs, guiding action selection and resulting in a feedback loop where actions bring sensory data closer to expectations. It is important to note difficulty in differentiation between the active and the perceptive, which brings us to the question of modeling collective desire: is there a way to tell the difference between when we act according to an “inner” logic we have conscious cognitive access to, and when are we moved by a logic we simply ‘feel’ driven by? Friston and others postulate that free energy minimization, as a principle, might reflect broader systemic efficiencies in biological scaling laws, where living systems seek to maintain optimal internal states (allostatic setpoints) with minimal energetic cost [Vasil et al., 2020]. The “average” aspect is important to emphasize here: for complex systems, long-term averages may signify significant shorter-term high-risk, high-energy investments.
@@
where $\ln\tilde{p}(o)$ represents the agent’s preferences over observations or extrinsic value, expressed as a (biased) prior probability distribution over observations [Parr et al., 2022]. The first term represents the (negative) expected information gain or epistemic value, which scores the degree to which an agent expects to update their beliefs in light of new sensory data. By pursuing outcomes that maximize a balance between epistemic and extrinsic values, agents simultaneously satisfy immediate goals and resolve uncertainty, driving adaptive behavior in complex environments.
@@ 1.2 Why active inference is suitable for modeling power dynamics
Active Inference Framework (AIF) is well-suited to understand power dynamics, beginning with individual agents as predictive systems that minimize uncertainty through perception and action [Bruineberg, 2017]. These agents develop control through embodied interaction with their environment, establishing what Bruineberg calls the “primacy of the ’I Can”’—the fundamental ability, or power, to influence one’s surroundings. This individual control scales up through generative models that enable agents to predict and understand others’ intentions [Friston et al., 2020, Vasil et al., 2020], while attention is fundamentally shaped by deep preferences that determine what becomes salient in the environment [Parvizi-Wayne, 2024].
@@
At the social level, power dynamics emerge through shared predictive models and environmental mediation. Groups develop shared narratives that coordinate actions and expectations [Bouizegarene et al., 2020], while stigmergic processes allow for distributed control through environmental modifications that influence collective behavior [Friedman et al., 2021]. This social coordination is reinforced through deontic cues that signal social norms and expectations [Constant et al., 2019], while cultural practices shape attention patterns to create shared perceptual landscapes within the environment. Joint attention and shared fields of affordances enable efficient coordination through mutual understanding of action possibilities [Tison and Poirier, 2021, Vasil et al., 2020].
@@
These mechanisms ultimately can give rise to complex power structures through decentralized interactions rather than top-down control [Friedman et al., 2021]. Groups self-organize into subgroups based on belief similarity [Kastel et al., 2023], while influence spreads through network structures that amplify certain messages while marginalizing others [Albarracin et al., 2022]. This process creates systems of ‘cultural capital’ [Bourdieu, 2018] where shared knowledge and norms enable efficient navigation of social contexts [Constant et al., 2019, Hipólito and van Es, 2022]. The result is a dynamic system where power emerges from the interplay of individual prediction, social influence, and cultural learning, all of which can be unified under the principles of AIF.
@@
The question of social sameness (as in “same” narratives, belief similarity, etc.) and (dis)attention to it, is therefore a political question. If attention is fundamentally shaped by deep preferences that determine what becomes salient in the environment for an agent or group, then we ought to ask how the production and performance of scripts (habitus, narratives: histories, identities, etc.) can be newly understood so that we may critically construct these collectively, without risking that this collectivity is induced by strong homogenizing scripts which occlude or incapacitate ever-renewing possibilities (as warned by Arendt). Incentives to promote legibility in a population can also manifest in institutions shaping and direction of attention toward specific narratives [Scott, 2020], and may similarly arise in collectively determined scripts. Thus, a balance needs to be maintained between benefits such as improved coordination and an oversuppression of diversity among constituent members of collectives. As mentioned earlier, “There is no alternative” can be understood as the poster-child of ‘one-dimensional’ [^14] For this Marcusian notion at work in similar ways in the domain of NLP, see [de Jager, 2023]. visions effectuated by reducing complex systems into simplified structures which pool attention towards a specific goal. Through our treatment, a new definition of social power can thus be understood as possibilistic. Essentially: if we ought to adopt a politics based on the knowledge that scripts work in the ways discussed above and in what follows below, and if the organization of life is hierarchical by definition (both physically and metaphysically: we are given to a world of already-existing constraints which we organize around by means of concepts aimed at future possibilities), taking account of this by understanding the nature of distributed generative models—over time: as generations, as the representation of historical narratives, etc.—offers a vision of possibilistic landscapes based on collective dynamics (and their analyzable mechanics) which can be framed beyond our current visions based on simplistic ideas of agency as individualized freedom and through atomized actors.
@@ 2.1 Scripts as frameworks for social behavior
Social scripts are blueprints that enable agents to navigate social situations and select situationally appropriate behaviors [Albarracin et al., 2021]. They represent both internalized cognitive schemas and externalized social orders that structure behavioral possibilities within cultural contexts. Scripts harness socially shared knowledge that prescribes normative standards for behavior, similar to how actors share a dramaturgical script. Strong scripts specify precise sequences of structured behavioral events that agents are expected to perform in certain social situations [Albarracin et al., 2021]. These scripts maintain causal and temporal relationships between events, enabling agents to make reliable inferences about others’ likely actions. For example, the North American restaurant script prescribes specific ordered sequences: being seated by the host, reviewing menus, ordering drinks, ordering food, eating, and paying the bill including tip [Albarracin et al., 2021].
@@
Weak scripts specify the typical features and semantic associations that agents encounter in particular event types, without strictly prescribing event order [Albarracin et al., 2021]. These scripts function as clusters of conceptually related possibilities that help agents interpret situations and identify appropriate actions. The semantic connections in weak scripts adapt agents’ perceptual fields by making certain action possibilities more salient based on contextual goals.
@@
Scripts help maintain power structures. Scripts normalize certain behavioral patterns by providing templates for ”appropriate” action, making deviations appear abnormal or inappropriate (Constant et al., 2018). This normalization reinforces existing power relations by making them appear natural and inevitable. They establish reliable sequences of interaction in favor of a current norm: whether agents following this are (critical) aware of the norm, or simply perform it, is often the question. They thus reduce uncertainty and create predictable social patterns that benefit those already in power [Albarracin et al., 2021]. These patterns become self-reinforcing as agents learn to expect and reproduce them. Given this predictability, scripts assign and maintain social roles by specifying appropriate behaviors for different agents based on their position within power structures [] (Constant et al., 2018). This role differentiation helps reproduce hierarchical relationships.
@@
Looking at scripts under AIF, they can be understood as shared generative models that enable coordinated social behavior. Scripts represent shared knowledge about the causal structure of social situations, allowing agents to make predictions about others’ behavior and select appropriate actions. These shared models facilitate social coordination by reducing uncertainty about interaction patterns. Scripts incorporate deontic cues that trigger specific behavioral policies, helping agents select appropriate actions in response to environmental signals [Constant et al., 2019]. These cues guide action selection by making certain policies more likely in specific contexts. Scripts are transmitted across generations through cultural learning processes, enabling the reproduction of social structures. This transmission occurs through both explicit instruction and implicit learning from environmental regularities.